The length of a rectangle is six times its width. If the area of the rectangle is 486 cm^2 , find its perimeter.
Answer by checkley79(3341)

L=6W
A=L*W
486=6W*W
486=6W^2
W^2=486/6
W^2=81
W=SQRT81
W=9 cm ANS. FOR THE WIDTH.
L=6*9=54 cm ANS. FOR THE LENGTH.
PROOF
9*54=486
496=496
P=2L+2W
P=2*54+2*9
P=108+18=126 cm ANS. FOR THE PERIMETER.